2 6. 2pts) Suppose a consumer advocacy group would like to conduct a survey to find the proportion p of consumers who bought the newest generation of an MP3 player who were happy with their purchase. How large a sample should they take to estimate p with 2% margin oferror and 90% confidence?
Answer #1

Z2/2 × p × (1-p) MOE2 rl

\alpha = 0.10

From Table, critical values of Z = \pm 1.645

Since p is not given:
Take p = 0.5

MOE = 0.02

Substituting, we get:

n=\frac{1.645^{2}\times 0.5\times (1-0.5)}{0.02^{2}}=1692

So,

Answer is:

1692
Answer #2

How can we just take p=0.5??

source: IIM
answered by: Ash